1. Maria can read 15 pages of economics in an hour. She can also read 40 pages of sociology in an hour. She spends 8 hours per day studying.

a. Draw Maria’s production possibilities frontier for reading economics and sociology.

b. What is Maria’s opportunity cost of reading 80 pages of sociology?


1
Expert's answer
2022-04-04T09:15:01-0400

2.To read 1 page of Sociology, she is giving up,20pages


50pages

​=2.5 pages of EconomicsTherefore, to read 100 pages of Economics,Maria would have to give up(100∗2.5)=250pages of Economicswhich is her opportunity cost.

"\text{1.}\, \text{E}=20\ \text{pages}\\

\text{S}=50\ \text{pages}\\

\text{If Maria spends 5hours on reading Economics}\\

\text{she would read} (20*5)\text{pages}=100\text{pages}\\

\text{If she spends 5hours on Sociology,}\\

\text{she would read} (50*5)\text{pages}=250\text{pages}\\

\text{if she decides to spend 2.5hours each on them,}\\

\text{she would read};\\

(20*2.5)\text{pages}=50\text{pages of Economics}\\

(50*2.5)\text{pages}=125\text{pages of Sociology}\\"